Two dice are used, each die loaded so that the probabilities of 1,2,3,4,5,6 are x1-x/6, 1+2x/6, x-1/6, 1+x/6, 1-2x/6, 1+x/6 respectively.

Compute the probability in rolling the two dice sum up to six


1
Expert's answer
2021-08-12T18:18:44-0400

Solution:

Using "\Sigma p_i=1"

"(\dfrac{x}{1}-\dfrac{x}{6})+(\dfrac{1+2x}{6})+(x-\dfrac{1}{6})+(\dfrac{1+x}{6})+(\dfrac{1-2x}{6})+(\dfrac{1+x}{6})=1
\\\Rightarrow x=\dfrac3{13}"

Then, the probability distribution:

"X=1,p=\dfrac{5}{26}
\\ X=2,p=\dfrac{19}{78}
\\ X=3,p=\dfrac{5}{78}
\\ X=4,p=\dfrac{8}{39}
\\ X=5,p=\dfrac{7}{78}
\\ X=6,p=\dfrac{8}{39}"

"P(S\le 6)=P(X=1,X=1)+P(X=1,X=2)+...+P(X=1,X=5)
\\+P(X=2,X=1)+P(X=2,X=2)+...+P(X=2,X=4)
\\+P(X=3,X=1)+P(X=3,X=2)+...+P(X=3,X=3)
\\+P(X=4,X=1)+P(X=4,X=2)
\\+P(X=5,X=1)"

"=\dfrac{5}{26}\times \dfrac{5}{26}+\dfrac{5}{26}\times \dfrac{19}{78}+\dfrac{5}{26}\times\dfrac{5}{78}+\dfrac{5}{26}\times\dfrac{8}{39}+\dfrac{5}{26}\times\dfrac{7}{78}"

"+\dfrac{19}{78}\times \dfrac{5}{26}+\dfrac{19}{78}\times \dfrac{19}{78}+\dfrac{19}{78}\times \dfrac{5}{78}+\dfrac{19}{78}\times \dfrac{8}{39}"

"+\dfrac{5}{78}\times\dfrac{5}{26} +\dfrac{5}{78}\times\dfrac{19}{78} +\dfrac{5}{78}\times \dfrac{5}{78}"

"+\dfrac{8}{39}\times \dfrac{5}{26}+\dfrac{8}{39}\times\dfrac{19}{78}" "+\dfrac{7}{78}\times\dfrac{5}{26}"

"=\dfrac{155}{1014}+\dfrac{1045}{6084}+\dfrac{5}{156}+\dfrac{136}{1521}+\dfrac{35}{2028}
\\=\dfrac{2819}{6084}"